Meeting Notes for 10 May

Present: Sudha, Marge, Mark, Ken, Tom, Joel


  • Ken has tested the site features extensively, with notes posted to BaseCamp. He'll focus on testing Groups this week.
  • Sudha will be able to take care of several of Ken's notes herself. She'll get Xeno Media's input on the rest.


No specific updates this week. Rolling out updated code to the e-Lab is delayed to the mid-May / early June range. Ken has received many inquiries about CMS summer workshops, so updating the e-Lab remains a priority despite the delay.


  • Mark and Nathan Unterman are working on detector hardware for the August eclipse project.
  • Mark will review CREDO's work with the sample data later in the week. Based on his assessment, Joel will begin transferring data. They've been allocated another 20TB of space, which will be more than sufficient.
  • Sudha has created an "eclipse data" file out of the Shower Analysis "eventCandidates" file. She's now working on the data conversion code and its integration into the e-Lab Analysis system.
  • Ken has had persistent problems with network connection resets while uploading large CR data sets. It may be a local network problem. He'll fill out a Help Desk Request to see if other users have encountered anything similar.


The VMs and physical servers will be down for at least some of the 8AM - 11AM block on Saturday, May 20 for annual CRC maintenance. They will be upgrading their own network infrastructure as well as patching our physical servers. Joel will patch the VMs concurrently. This may interfere with the activities of the Fellows' Workshop at Fermilab. Marge and Ken will see what might be affected, and Joel will check with the CRC about how flexible the schedule can be.

-- Main.JoelG - 2017-05-10


Topic revision: r1 - 2017-05-10, JoelG
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback